OpenStudy (anonymous):

find the limit if it exists: lim Δx->0 ((x+Δx)^2-x^2)/(Δx)

OpenStudy (anonymous):

i tried factoring out the swuared part but i messed up somewhere, im taking another shot at it

OpenStudy (anonymous):

Start by expanding \[(x + Deltax)^{2}\]

OpenStudy (anonymous):

(x+Δx)(x+Δx)=x^2+Δx^2+Δx^2+2Δx^2-x^2/Δx. that is what i got. i think it is wrong though.

OpenStudy (anonymous):

\[(x + \Delta x)^{2}\] = \[x ^{2} + (\Delta x)^{2} + 2x(\Delta x)\]

OpenStudy (anonymous):

the delta x is just a single variable. It looks confusing because it requires more than one symbol.

OpenStudy (anonymous):

im going to try again

OpenStudy (anonymous):

so the first step was to get: x^2+x(Δx)+x(Δx)+(Δx)^2?

OpenStudy (anonymous):

Once you subtract out an x^2 from what I showed you, you'll have 2 terms left that you can factor the delta x out of because of the delta x in the denominator.

OpenStudy (anonymous):

by doing that i get 2x(Δx)+Δx^2 / Δx

OpenStudy (anonymous):

Yes, that was the first part but not the key. The key is to get rid of the delta x in the denominator.

OpenStudy (anonymous):

yes i know. im just taking it step by step. this is my first time seeing this delta symbol being used in this manner

OpenStudy (anonymous):

np, you're doing well. So, the numerator can factor out a delta x.

OpenStudy (anonymous):

And because you'll have a factor of delta x in both the numerator and denominator, those factors will cancel.

OpenStudy (anonymous):

hey, im going to type something out. tell me if this is true

OpenStudy (anonymous):

i tried factoring out the numerator and got this. (Δx)(2x+Δx)/(Δx). so i am left with just 2x+Δx.

OpenStudy (anonymous):

Very, very good. That's the key. Now there's no denominator to worry about as delta x goes to 0.

OpenStudy (anonymous):

yay~!

OpenStudy (anonymous):

yay is key! Actually, groovy!

OpenStudy (anonymous):

you cleared things up big time by having me look at (Δx) as one thing.

OpenStudy (anonymous):

Math is way cool when the lightbulb goes on.

OpenStudy (anonymous):

yep. so now 0 is plugged in and i get 2(0)+Δ(0)=0

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

you are left with 2x + delta x. It is only the delta x that is going to 0 so you are left with 2x once delta x goes to 0

OpenStudy (anonymous):

oh wait the lmiit is Δx. so only 0 can replace Δx, correct?

OpenStudy (anonymous):

i get ahead of myself sometimes :\

OpenStudy (anonymous):

Yes, only delta x is approaching 0. x stays the same.

OpenStudy (anonymous):

ahh i see. so the answer would be lim Δx-> = 2x

OpenStudy (anonymous):

That's it! and with that, you have the answer!

OpenStudy (anonymous):

aweeeeeesome. thank you so much!

OpenStudy (anonymous):

thx for the medal and more importantly, good luck to you always in all your studies. Learning can be very fun.

OpenStudy (anonymous):

yea, sure is!